OpenStudy (anonymous):

Solve t = p+3r/6 for r.

OpenStudy (anonymous):

Show all work... Im lost

OpenStudy (mimi_x3):

Omg ! worked it out but not sure if I'm right though xD I got r = p/3 + 3t , is that right ?

OpenStudy (anonymous):

Im not sure...

OpenStudy (mimi_x3):

w8 , -p/3 + 3t Don't you have the answers to it ? xD

OpenStudy (anonymous):

2t-2p=r

OpenStudy (anonymous):

its a question for homework... which i am completely lost with. the practice problems in school were completly different then now.

OpenStudy (anonymous):

6t-p/3=r?

OpenStudy (anonymous):

Is the equation t = p+ (3r/6) or is it t = (p+3r)/6. If it is the first one, I got 2(t-p) = r. It is all a manner of just moving everything around.

OpenStudy (anonymous):